# SCMM: Calibrating Cross-modal Representations for Text-Based Person Search

### v7

@@ 3.1 Sew Calibration Loss with Constraints
(a) Single-modal constraints
(b) Cross-modal constraints
@@
Subsequently, we change the above pairwise constraints into soft forms for better convergence following [33].
@@
where $\bar{v}^{c}$ and $\bar{t}^{c}$ are CLS token features after normalizing, $K$ and $J$ denote the number of positive and negative samples in this batch, respectively. $\alpha$ is a scale parameter. As a result, the image-to-text matching part of sew calibration loss is formulated as below:
@@
The constraints in Eq. 2 can also be used to impose a classification loss in a similar way. As there is no difference for perfect positive samples in the classification task, we omit to constrain Eq. 1. Formally, the loss for our person ID classification part is as follows [16]:
$\mathcal{L}^{I2T}_{id}&=\frac{1}{n}\sum\limits_{i=1}^{n}-log\big{(}\frac{e^{(\alpha(s_{y_{i},i}-\mathcal{M}_{2}))}}{e^{(\alpha(s_{y_{i},i}-\mathcal{M}_{2}))}+\sum\limits_{j\neq{y_{i}}}e^{(\alpha(s_{y_{j},i}))}}),\end{aligned}$ (8)
@@
3.2 Masked Caption Modeling Loss
TBPS is a fine-grained cross-modal task, which means only caption-level discrimination is not enough. If the textual captions of two persons differ in a few words, a TBPS method can not retrieve a specific person without word-level discrimination. Although there are many works to establish word-level discrimination capacities, such methods are complex and limited, hindering performance boost. To solve this issue, we propose masked caption modeling to establish detailed image-text relationships. Furthermore, by utilizing MCM, our framework can perform more generic cross-modal learning.
Inspired by [6, 35], we add a masked prediction task on the text branch. Concretely, this loss is based on a cross-modal decoder architecture. We mask a portion of text tokens and replace these masked tokens with a learnable token vector. The text encoder inputs these text tokens and outputs the corresponding text features. The cross-modal decoder $f_{cd}$ learns to maximize the conditional likelihood of the masked text feature $t_{n}$ under latent image feature sequence $\{v_{i}\}$ and text feature sequence $\{t_{i}\}$:
@@ 3.2 Masked Caption Modeling Loss
where $N$ is the total masked token numbers in a caption.
Input:
Image $I$ and text $T$; A batch of $n$ paired $\mathbb{G}=\{(I^{1},T^{1}),(I^{2},T^{2}),\cdots,(I^{n},T^{n})\}$
@@
foreach $T^{i}$ in set $\mathbb{G}$ do
      $T^{i}\leftarrow\text{mask}(T^{i})$;
end foreach
for $i\leftarrow 1$ to $\mathbb{G}$ do
       $(I^{i}_{cls},I^{i}_{1},\cdots,I^{i}_{c})\leftarrow ViT(I^{i})$;
       $(T^{i}_{cls},T^{i}_{1},\cdots,T^{i}_{c})\leftarrow Bert(T^{i})$;
       if Training Stage then
             $(I^{i}_{cls},T^{i}_{cls})\leftarrow BatchNorm(I^{i}_{cls},T^{i}_{cls})$;
             $I^{i}_{context},T^{i}_{context}\leftarrow(I^{i}_{1},\cdots,I^{i}_{c}),(T^{i}_{1},\cdots,T^{i}_{c})$;
             $T^{i}_{context}\leftarrow Attn(T^{i}_{context})$;
             $T^{i}_{cross}\leftarrow CrossAttn(T^{i}_{context},I^{i}_{context})$;
             $\mathcal{L}_{sew}^{i}\leftarrow SewCalibration(I^{i}_{cls},T^{i}_{cls})$;
             $\mathcal{L}_{mcm}^{i}\leftarrow MCM(T^{i}_{cross},T^{i})$;
      else
             $\{(I^{1}_{cls},T^{1}_{cls}),(I^{2}_{cls},T^{2}_{cls}),\cdots,(I^{i}_{cls},T^{i}_{cls})\}\leftarrow(I_{cls},T_{cls})$;
       end if
end for
if Training Stage then
       return $(\mathcal{L}_{sew},\mathcal{L}_{mcm})$;
else
       return $(I_{cls},T_{cls})$;
end if
@@
As shown in Fig. 2, the cross-modal decoder part, the decoder $f_{cd}$ takes both unmasked text tokens and masked tokens in their original order as the input. The multi-head self-attention [18] first encodes the text features as $Q_{t},K_{t},V_{t}$, while the cross self-attention further improves the text features by taking into account the encoded image features as $K_{i},V_{i}$ for visual context. The final linear projection layer has the same number of output channels as the text vocabulary and computes the cross entropy loss between the reconstructed and original words only on masked text tokens. It should be noted that $f_{cd}$ is only used during training and not during inference.
3.3 Total Loss
The total loss function of SCMM integrates our two complementary fusion mechanisms to achieve comprehensive cross-modal alignment and correspondence learning. Our unified optimization objective combines the sew calibration loss and the masked caption modeling loss to jointly address both global alignment and fine-grained correspondence challenges in cross-modal information fusion. The sew calibration loss $\mathcal{L}_{sew}$ encompasses bidirectional matching and classification components that ensure symmetric cross-modal alignment:
@@ 3.3 Total Loss
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters balancing global alignment and fine-grained correspondence learning. Through this joint optimization, the framework simultaneously learns compact cross-modal representations while capturing detailed visual-textual relationships for discriminating similar individuals. The complementary loss components create synergistic effects ensuring comprehensive information fusion, addressing both inter-modal gap reduction and fine-grained correspondence modeling. Regarding computational complexity, SCMM maintains efficient processing through its streamlined dual-encoder architecture. Training complexity is $O(n\cdot d^{2})$ where $n$ is batch size and $d$ is feature dimension, dominated by cross-attention operations in the MCM decoder. During inference, only CLS token embeddings are utilized, achieving $O(d)$ complexity per query-gallery comparison. In contrast, the method contrasts favorably with attention-explicit methods like ACSA [26] requiring $O(n^{2}\cdot d)$ complexity due to full cross-modal attention during inference. Our training-only decoder design ensures high-speed retrieval while maintaining sophisticated cross-modal learning capabilities.
4 Experiments
4.1 Datasets and Evaluation Metric
@@
We evaluate SCMM on three benchmarks: CUHK-PEDES [8], ICFG-PEDES [20], and RSTPReid [13]. CUHK-PEDES is the first large-scale benchmark for text-based person search tasks. As the foundational dataset, it contains 40,206 images of 13,003 person IDs collected from five person re-identification datasets. Each image has two different textual captions with an extensive vocabulary, and the average sentence length is 23.5. The testing set comprises 3,074 images and 6,148 descriptions of 1,000 persons. ICFG-PEDES contains 54,522 images of 4,102 persons. For each image, the corresponding description has an average length of 37 words. The testing set consists of 19,848 image-text pairs of 1,000 persons. RSTPReID contains 20,505 images of 4,101 persons, with each pedestrian having five images. It is divided into a training set with 3,701 persons, a validation set with 200 persons, and a testing set with 200 persons. For our evaluation metric, we report the Rank $k$ ($k$=1, 5, 10) text-to-image accuracy, which is commonly used in previous works to evaluate text-based person search. Given a textual description as the query, if the top-$k$ retrieved images contain any person corresponding to the query, we consider it a successful person search.
4.2 Implementation Details
@@
In our visual-textual dual-encoder, we extract visual representations using the ViT-Base pre-trained on ImageNet [38]. The images are resized to 224 $\times$ 224 pixels. For textual representations, we use the BERT-Base-Uncased model pre-trained on the Toronto Book Corpus and Wikipedia. The representation dimension is set to 768, and the feature sequence lengths are set to 197 and 100, respectively. During the training phase, we use a batch size of 64 and train for 60 epochs. We use Adam optimizer with an initial learning rate of 0.001. To augment our data, we apply a random horizontal flipping operation, and we use a mask ratio of 0.1 for randomly masking text tokens. The minimum and maximum textual information length boundaries $\mathcal{T}_{min}$ and $\mathcal{T}_{max}$ are set to 20-60, 25-65, and 22-60 according to the caption length distributions of CUHK-PEDES, ICFG-PEDES, and RSTPReID, respectively. The bounds of the upper and lower margin $\mathcal{M}$ are set to 0.4 and 0.6, and the scale parameter $\alpha$ is set to 32. For each loss in the total loss function, the balance factors $\lambda_{1}$ and $\lambda_{2}$ are set equal to 1. During the testing phase, we apply a re-ranking post-processing approach to improve search performance following NAFS [14]. We conduct the experiments on four NVIDIA 2080Ti GPUs using PyTorch.
4.3 Comparison with State-of-the-art Methods
@@
Table 1: Comparison with SOTA methods on the CUHK-PEDES dataset. Rank1, Rank5, and Rank10 accuracies (%) are reported. Bold number represents the best score. R and C denote the re-ranking post-processing operations and image encoder pre-trained on CLIP model.
@@
GNA-RNN [8] 19.05 - 53.64
CMPM/C [16] 49.37 - 79.27
Dual-path [15] 44.40 66.26 75.07
ViTAA [19] 55.97 75.84 83.52
VP Net [27] 58.83 81.25 86.72
SUM [39] 59.22 80.35 87.60
CLIP [28] 60.67 81.99 88.87
DSSL [13] 59.98 80.41 87.56
MGEL [23] 60.27 80.01 86.74
SSAN [20] 61.37 80.15 86.73
TestReID [40] 64.08 81.73 88.19
ACSA [26] 63.56 81.40 87.70
IVT [37] 64.00 82.72 88.95
SAFA Net [22] 64.13 82.62 88.40
TIPCB [21] 64.26 83.19 89.10
CAIBC [12] 64.43 82.87 88.37
AXM Net [41] 64.44 80.52 86.77
CFine [25] 65.07 83.01 89.00
ISANet [5] 63.92 82.15 87.69
@@
4.3.1 Results on CUHK-PEDES.
Table 1 compares our framework and previous methods on CUHK-PEDES. It can be observed that our proposed SCMM can outperform all previous methods by a large margin. Compared to the work CFine [25], SCMM achieves 67.71% (+2.64%), 84.57% (+1.56%) and 89.44% (+0.44%) on Rank1, Rank5 and Rank10 without re-ranking. The state-of-the-art results on the CUHK-PEDES show the effectiveness of SCMM. With the help of re-ranking, SCMM shows an incremental boost to get a 71.09% Rank1 score. Furthermore, with a better image encoder pre-training on CLIP, SCMM achieves 73.81%, 88.89% , and 92.77% on three metrics, respectively. Overall, the consistent improvements show the scalability of SCMM across better pre-trained models and extra post-processing operations.
Table 2: Quantitative results on the ICFG-PEDES dataset.
@@
CMPM/C [16] 43.51 65.44 74.26
Dual-path [15] 38.99 59.44 68.41
ViTAA [19] 50.98 68.79 75.78
SSAN [20] 54.23 72.63 79.53
CLIP [28] 53.96 73.69 80.43
TIPCB [21] 54.96 74.72 81.89
IVT [24] 56.04 73.60 80.22
CFine [25] 55.69 72.72 79.46
CFine+C [25] 60.83 76.55 82.42
IRRA [34] 63.46 80.25 85.82
TP-TPS [11] 60.64 75.97 81.76
ISANet [5] 57.73 75.42 81.72
CANC [2] 60.52 78.36 84.13
@@
Table 3: Quantitative results on the RSTPReid dataset.
@@
CLIP [28] 50.10 76.10 84.95
DSSL [13] 32.43 55.08 63.19
SUM [39] 41.38 67.48 76.48
SSAN [20] 43.50 67.80 77.15
IVT [24] 46.70 70.00 78.80
ACSA [26] 48.40 71.85 81.45
CFine [25] 45.85 70.30 78.40
CFine+C [25] 50.55 72.50 81.60
TP-TPS [11] 50.65 72.45 81.20
@@
4.3.2 Results on ICFG-PEDES and RSTPReid.
We also utilized other benchmarks to validate SCMM’s performance and generalization. The ICFG-PEDES and RSTPReid datasets are more challenging compared to CUHK-PEDES, and our method significantly outperformed all state-of-the-art methods on these two datasets by a large margin, as reported in Tables 2 and 3. Compared with the state-of-the-art results [25] on ICFG-PEDES, SCMM achieved significant improvements, with scores of 60.20% (+4.51%), 75.97% (+3.25%), and 81.78% (+2.32%) on Rank1, Rank5, and Rank10, respectively. On the RSTPReid dataset, we achieved scores of 50.75% (+4.90%), 74.20% (+3.90%), and 81.70% (+3.30%) on the three metrics. With re-ranking post-processing, we were able to achieve scores of 62.17% and 55.35% on Rank1 for the two benchmarks, respectively.
@@
4.4 Ablation Studies
Table 4: Performance comparisons of different components. Res and ViT denote ResNet and Vision Transformer as image encoders, respectively. Sew-F and Sew-A refer to the fixed and adaptive margin versions of the sew calibration loss, respectively. Model 2 is our baseline.

### v8

@@ III-A Sew Calibration Loss with Constraints
Subsequently, we change the above pairwise constraints into soft forms for better convergence following [circle].
@@
where $\bar{v}^{c}$ and $\bar{t}^{c}$ are CLS token features after normalizing, $K$ and $J$ denote the number of positive and negative samples in this batch, respectively. $\alpha$ is a scale parameter. As a result, the image-to-text matching part of the sew calibration loss is formulated as below:
@@
The constraints in Eq. 2 can also be used to impose a classification loss in a similar way. As there is no difference for perfect positive samples in the classification task, we omit to constrain Eq. 1. Formally, the loss for our person ID classification part is as follows [cmpmcmpc]:
$\begin{aligned} \mathcal{L}^{I2T}_{id}&=\frac{1}{n}\sum\limits_{i=1}^{n}-log\big(\frac{e^{(\alpha(s_{y_{i},i}-\mathcal{M}_{2}))}}{e^{(\alpha(s_{y_{i},i}-\mathcal{M}_{2}))}+\sum\limits_{j\neq{y_{i}}}e^{(\alpha(s_{y_{j},i}))}}),\end{aligned}$ (8)
@@
III-B Masked Caption Modeling Loss
TBPS is a fine-grained cross-modal task, which means that only caption-level discrimination is not enough. If the textual captions of two persons differ in a few words, a TBPS method can not retrieve a specific person without word-level discrimination. Although there are many works to establish word-level discrimination capacities, such methods are complex and limited, hindering performance boost. To solve this issue, we propose masked caption modeling to establish detailed image-text relationships. Furthermore, by utilizing MCM, our framework can perform more generic cross-modal learning.
Inspired by [zhai2022lit, devlin2018bert], we add a masked prediction task on the text branch. Concretely, this loss is based on a cross-modal decoder architecture. We mask a portion of text tokens and replace these masked tokens with a learnable token vector. The text encoder inputs these text tokens and outputs the corresponding text features. The cross-modal decoder $f_{cd}$ learns to maximize the conditional likelihood of the masked text feature $t_{n}$ under latent image feature sequence $\{v_{i}\}$ and text feature sequence $\{t_{i}\}$:
@@ III-B Masked Caption Modeling Loss
where $N$ is the total masked tokens numbers in a caption.
Input: Image II and text $T$; A batch of $n$ paired $\mathbb{G}=\{(I^{1},T^{1}),(I^{2},T^{2}),\cdots,(I^{n},T^{n})\}$
@@
1 foreach $T^{i}$ in set $\mathbb{G}$ do
2 $T^{i}\leftarrow\text{mask}(T^{i})$;
3 end foreach
4for $i\leftarrow 1$ to $\mathbb{G}$ do
5 $(I^{i}_{cls},I^{i}_{1},\cdots,I^{i}_{c})\leftarrow ViT(I^{i})$;
6 $(T^{i}_{cls},T^{i}_{1},\cdots,T^{i}_{c})\leftarrow Bert(T^{i})$;
7 if Training Stage then
8    $(I^{i}_{cls},T^{i}_{cls})\leftarrow BatchNorm(I^{i}_{cls},T^{i}_{cls})$;
9    $I^{i}_{context},T^{i}_{context}\leftarrow(I^{i}_{1},\cdots,I^{i}_{c}),(T^{i}_{1},\cdots,T^{i}_{c})$;
10    $T^{i}_{context}\leftarrow Attn(T^{i}_{context})$;
11    $T^{i}_{cross}\leftarrow CrossAttn(T^{i}_{context},I^{i}_{context})$;
12    $\mathcal{L}_{sew}^{i}\leftarrow SewCalibration(I^{i}_{cls},T^{i}_{cls})$;
13    $\mathcal{L}_{mcm}^{i}\leftarrow MCM(T^{i}_{cross},T^{i})$;
14
15 else
16    $\{(I^{1}_{cls},T^{1}_{cls}),(I^{2}_{cls},T^{2}_{cls}),\cdots,(I^{i}_{cls},T^{i}_{cls})\}\leftarrow(I_{cls},T_{cls})$;
17
18   end if
19
20 end for
21if Training Stage then
22   return $(\mathcal{L}_{sew},\mathcal{L}_{mcm})$;
23
24else
25   return $(I_{cls},T_{cls})$;
26
27 end if
@@ 13    $\mathcal{L}_{mcm}^{i}\leftarrow MCM(T^{i}_{cross},T^{i})$;
As shown in Fig. 2, the cross-modal decoder part, the decoder $f_{cd}$ takes both unmasked text tokens and masked tokens in their original order as the input. The multi-head self-attention [attention] first encodes the text features as $Q_{t},K_{t},V_{t}$, while the cross self-attention further improves the text features by taking into account the encoded image features as $K_{i},V_{i}$ for visual context. The final linear projection layer has the same number of output channels as the text vocabulary and computes the cross-entropy loss between the reconstructed and original words only on masked text tokens. It should be noted that $f_{cd}$ is only used during training and not during inference.
III-C Total Loss
The total loss function of SCMM integrates our two complementary learning mechanisms to achieve comprehensive cross-modal alignment and correspondence learning. Our unified optimization objective combines the sew calibration loss and the masked caption modeling loss to jointly address both global alignment and fine-grained correspondence challenges in cross-modal representation learning. The sew calibration loss $\mathcal{L}_{sew}$ includes bidirectional matching and classification components that ensure symmetric cross-modal alignment:
@@ III-C Total Loss
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters balancing global alignment and fine-grained correspondence learning. Through this joint optimization, the framework simultaneously learns compact cross-modal representations while capturing detailed visual-textual relationships for discriminating similar individuals. The complementary loss components create synergistic effects ensuring comprehensive representation learning, addressing both inter-modal gap reduction and fine-grained correspondence modeling. Regarding computational complexity, SCMM maintains efficient processing through its streamlined dual-encoder architecture. Training complexity is $O(n\cdot d^{2})$ where $n$ is the batch size and $d$ is the feature dimension, dominated by cross-attention operations in the MCM decoder. During inference, only CLS token embeddings are utilized, achieving $O(d)$ complexity per query-gallery comparison. In contrast, the method contrasts favorably with attention-explicit methods like ACSA [acsa] requiring $O(n^{2}\cdot d)$ complexity due to full cross-modal attention during inference. Our training-only decoder design ensures efficient retrieval while maintaining sophisticated cross-modal learning capabilities. The pseudocode of our framework pipeline is shown in Alg. 1.
IV Experiments
IV-A Datasets and Evaluation Metric
@@
We evaluate SCMM on three benchmarks: CUHK-PEDES [gnarnn], ICFG-PEDES [ssan], and RSTPReid [dssl]. As the first large-scale benchmark for text-based person search tasks, CUHK-PEDES contains 40,206 images of 13,003 person IDs collected from five person re-identification datasets, with each image having two different textual captions with an extensive vocabulary and an average sentence length of 23.5 words. Additionally, the testing set comprises 3,074 images and 6,148 descriptions of 1,000 persons. ICFG-PEDES contains 54,522 images of 4,102 persons, where each image’s corresponding description has an average length of 37 words, and the testing set consists of 19,848 image-text pairs of 1,000 persons. RSTPReID contains 20,505 images of 4,101 persons, with each pedestrian having five images. It is divided into a training set with 3,701 persons, a validation set with 200 persons, and a testing set with 200 persons. For our evaluation metric, we report the Rank $k$ (kk=1, 5, 10) text-to-image accuracy, which is commonly used in previous works to evaluate text-based person search. Given a textual description as the query, if the top-$k$ retrieved images contain any person corresponding to the query, we consider it a successful person search.
IV-B Implementation Details
@@
In our visual-textual dual-encoder, we extract visual representations using the ViT-Base pre-trained on ImageNet [imagenet]. The images are resized to 224 $\times$ 224 pixels. For textual representations, we use the BERT-Base-Uncased model pre-trained on the Toronto Book Corpus and Wikipedia. The representation dimension is set to 768, and the feature sequence lengths are set to 197 and 100, respectively. During the training phase, we use a batch size of 64 and train for 60 epochs. We use Adam optimizer with an initial learning rate of 0.001. To augment our data, we apply a random horizontal flipping operation, and we use a mask ratio of 0.1 for randomly masking text tokens. The minimum and maximum textual information length boundaries $\mathcal{T}_{min}$ and $\mathcal{T}_{max}$ are set to 20-60, 25-65, and 22-60 according to the caption length distributions of CUHK-PEDES, ICFG-PEDES, and RSTPReID, respectively. The bounds of the upper and lower margin $\mathcal{M}$ are set to 0.4 and 0.6, and the scale parameter $\alpha$ is set to 32. For each loss in the total loss function, the balance factors $\lambda_{1}$ and $\lambda_{2}$ are set equal to 1. During the testing phase, we apply a re-ranking post-processing approach to improve search performance following NAFS [nafs]. We conduct the experiments on four NVIDIA 2080Ti with PyTorch.
IV-C Comparison with State-of-the-art Methods
@@
TABLE I: Comparison with SOTA methods on the CUHK-PEDES dataset. Rank1, Rank5, and Rank10 accuracies (%) are reported. Bold number represents the best score. R and C denote the re-ranking post-processing operations and image encoder pre-trained on CLIP model.
@@
GNA-RNN [gnarnn] 19.05 - 53.64
CMPM/C [cmpmcmpc] 49.37 - 79.27
Dual-path [dualpath] 44.40 66.26 75.07
ViTAA [vitaa] 55.97 75.84 83.52
VP Net [verbalNet] 58.83 81.25 86.72
SUM [sum] 59.22 80.35 87.60
CLIP [clip] 60.67 81.99 88.87
DSSL [dssl] 59.98 80.41 87.56
MGEL [mgel] 60.27 80.01 86.74
SSAN [ssan] 61.37 80.15 86.73
TestReID [limit] 64.08 81.73 88.19
ACSA [acsa] 63.56 81.40 87.70
IVT [vit] 64.00 82.72 88.95
SAFA Net [safaNet] 64.13 82.62 88.40
TIPCB [tipcb] 64.26 83.19 89.10
CAIBC [caibc] 64.43 82.87 88.37
AXM Net [axm] 64.44 80.52 86.77
CFine [cfine] 65.07 83.01 89.00
ISANet [isa] 63.92 82.15 87.69
CAPL [niu2024comprehensive] 65.63 84.81 90.21
LERF [zhang2025localenhanced] 65.84 84.24 90.22
@@
IV-C1 Results on CUHK-PEDES.
Table I compares our framework and previous methods on CUHK-PEDES. It can be observed that our proposed SCMM can outperform all previous methods by a large margin. Compared to the work CFine [cfine], SCMM achieves 67.71% (+2.64%), 84.57% (+1.56%) and 89.44% (+0.44%) on Rank1, Rank5 and Rank10 without re-ranking. The state-of-the-art results on the CUHK-PEDES show the effectiveness of SCMM. With the help of re-ranking, SCMM shows an incremental boost to get a 71.09% Rank1 score. Furthermore, with a better image encoder pre-training on CLIP, SCMM achieves 73.81%, 88.89% , and 92.77% on three metrics, respectively. Overall, the consistent improvements show the scalability of SCMM across better pre-trained models and extra post-processing operations.
TABLE II: Quantitative results on the ICFG-PEDES dataset.
@@
CMPM/C [cmpmcmpc] 43.51 65.44 74.26
Dual-path [dualpath] 38.99 59.44 68.41
ViTAA [vitaa] 50.98 68.79 75.78
SSAN [ssan] 54.23 72.63 79.53
CLIP [clip] 53.96 73.69 80.43
TIPCB [tipcb] 54.96 74.72 81.89
IVT [ivt] 56.04 73.60 80.22
CFine [cfine] 55.69 72.72 79.46
CFine+C [cfine] 60.83 76.55 82.42
IRRA [jiang2023crossmodal] 63.46 80.25 85.82
TP-TPS [bai2023textbased] 60.64 75.97 81.76
ISANet [isa] 57.73 75.42 81.72
LERF [zhang2025localenhanced] 57.23 76.64 83.11
CANC [gong2024crossmodal] 60.52 78.36 84.13
@@
TABLE III: Quantitative results on the RSTPReid dataset.
@@
  CLIP [clip]   50.10   76.10   84.95
  DSSL [dssl]   32.43   55.08   63.19
  SUM [sum]   41.38   67.48   76.48
  SSAN [ssan]   43.50   67.80   77.15
  IVT [ivt]   46.70   70.00   78.80
  ACSA [acsa]   48.40   71.85   81.45
  LERF [zhang2025localenhanced]   46.75   71.30   81.60
  CFine [cfine]   45.85   70.30   78.40
  CFine+C [cfine]   50.55   72.50   81.60
  TP-TPS [bai2023textbased]   50.65   72.45   81.20
@@
IV-C2 Results on ICFG-PEDES and RSTPReid.
We also utilized other benchmarks to validate SCMM’s performance and generalization. The ICFG-PEDES and RSTPReid datasets are more challenging compared to CUHK-PEDES, and our method significantly outperformed all state-of-the-art methods on these two datasets by a large margin, as reported in Tables II and III. Compared with the state-of-the-art results [cfine] on ICFG-PEDES, SCMM achieved significant improvements, with scores of 60.20% (+4.51%), 75.97% (+3.25%), and 81.78% (+2.32%) on Rank1, Rank5, and Rank10, respectively. On the RSTPReid dataset, we achieved scores of 50.75% (+4.90%), 74.20% (+3.90%), and 81.70% (+3.30%) on the three metrics. With re-ranking post-processing, we were able to achieve scores of 62.17% and 55.35% on Rank1 for the two benchmarks, respectively.
@@
IV-D Ablation Studies
TABLE IV: Performance comparisons of various components.